OpenStudy (anonymous):

A circle has an area of 36π square centimeters. What is the diameter?

OpenStudy (anonymous):

Area of circle=pi*radius^2 radius=1/2*diameter. 36pi=pi*radius^2 36=radius^2 radius=6 radius=1/2*diameter 6=1/2*diameter diameter=12 cm.

OpenStudy (anonymous):

area of circle = pi * r^2 = pi * (d/2)^2 since diameter/2 = raidus so 36 * pi sqin = pi * (d/2)^2 divide both sides by pi..... 36 *pi / pi = (d/2)^2 take the square root of both sides..... sqroot(36) = d/2 multiple both sides by 2 2* sqroot(36) =d d = 2*6 = 12 so the guy above is also right